# The Central Role of the Loss Function in Reinforcement Learning

### v2

@@ 4.4 First-Order Bounds for Online RL with Optimistic BCE Regression
Under Assump. 3, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 2 with the bce loss $\ell_{\operatorname{bce}}$ and $\beta=2\ln(H|\mathcal{F}|/\delta)$ enjoys the following:
$\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\widetilde{\mathcal{O}}(H\sqrt{{\color[rgb]{1,0,0}V^{\star}K}\cdot d\beta}+H^{2}d\beta),$
@@
We highlight that the only difference from Thm. 7 is changing the loss function to bce loss, so the more adaptive bound is truly a consequence of the loss function. Hence, this first-order bound can be specialized for low-rank MDPs by the same argument in Sec. 4.3. To the best of our knowledge, this is the first small-cost bound for low-rank MDPs in online RL without requiring distribution learning [61]. We now prove Thm. 8.
@@
Note that $\mathcal{T}_{h}^{\star}f_{h+1}(x_{h},a_{h})=\mathbb{E}[\tau^{\star}(f_{h+1},c_{h},x_{h+1})\mid x_{h},a_{h}]$, which is realizable by BC. Recall that $\delta_{h}^{\operatorname{Ber}}\leq\mathcal{E}_{h}^{\operatorname{bce}}$ by Lem. 2. We also write $\delta_{\operatorname{Ber}}^{\operatorname{RL}}=\sum_{h=1}^{H}\delta^{\operatorname{Ber}}_{h}$ and $\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{bce}}_{h}$. We now establish optimism.
Lemma 12.
Let $\ell=\ell_{\operatorname{bce}}$. Under the same setup as Lem. 7 with $\hat{f}^{\operatorname{op}}$ selected from $\mathcal{C}^{\operatorname{bce}}_{\beta}$ instead of $\mathcal{C}^{\operatorname{sq}}_{\beta}$, w.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}(\hat{f}^{\operatorname{op}},\pi^{i})\leq 2H\beta$, and (b) $\min_{a}\hat{f}^{\operatorname{op}}_{1}(x_{1},a)\leq V^{\star}$.
Proof of Lem. 12.
By Lem. 3 extended on martingale sequences [3], w.p.a.l. $1-\delta$, for all $f\in\mathcal{F}$, $h\in[H]$,
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{Ber}}_{h}(f,\pi^{i})$ $\textstyle\leq\ln(H|\mathcal{F}|/\delta)+\frac{1}{2}L_{h}^{\operatorname{bce}}(f_{h},f_{h+1},\mathcal{D}_{h})$
$\textstyle-\frac{1}{2}L_{h}^{\operatorname{bce}}(\mathcal{T}_{h}^{\star}f_{h+1},f_{h+1},\mathcal{D}_{h}).$ (18)
Let $g^{f}_{h}:=\mathop{\rm arg\,min}_{g_{h}\in\mathcal{G}_{h}}L_{h}(g_{h},f_{h+1},\mathcal{D}_{h})$ denote the empirical risk minimizer. Under the BC premise,
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{Ber}}_{h}(f,\pi^{i})$ $\textstyle\leq\ln(H|\mathcal{F}|/\delta)+\frac{1}{2}L_{h}^{\operatorname{bce}}(f_{h},f_{h+1},\mathcal{D}_{h})$
$\textstyle-\frac{1}{2}L_{h}^{\operatorname{bce}}(g^{f}_{h},f_{h+1},\mathcal{D}_{h}).$
Thus, any $f\in\mathcal{C}_{\beta}(\mathcal{D})$ satisfies $\sum_{i=1}^{n}\mathcal{E}^{\operatorname{Ber}}_{h}(f,\pi^{i})\leq\frac{1}{2}\beta+\ln(H|\mathcal{F}|/\delta)\leq 2\beta$, which proves Claim (a). For Claim (b), we prove that $Q^{\star}\in\mathcal{C}_{\beta}(\mathcal{D})$. By Eq. 18 and non-negativity of $\mathcal{E}^{\operatorname{Ber}}$, we have $L_{h}^{\operatorname{bce}}(\mathcal{T}_{h}^{\star}f_{h+1},f_{h+1},\mathcal{D}_{h})-L_{h}^{\operatorname{bce}}(g_{h}^{f},f_{h+1},\mathcal{D}_{h})\leq 2\ln(H|\mathcal{F}|/\delta)=\beta$. Then, setting $f=Q^{\star}$ and noting that $Q_{h}^{\star}=\mathcal{T}_{h}^{\star}Q^{\star}_{h+1}$ shows that $Q^{\star}$ satisfies the confidence set condition. Thus, $Q^{\star}\in\mathcal{C}_{\beta}(\mathcal{D})$ and Claim (b) follows by definition of $\hat{f}^{\operatorname{op}}$. ∎
By Lem. 12, we have $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\sum_{k=1}^{K}V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a)$. Then, the proof follows similarly as the squared loss case from before, except that we apply the finer Eq. 2 in place of Cauchy-Schwarz:
$\textstyle\sum_{k=1}^{K}V^{\pi^{k}}-f^{k}_{1}(x_{1},\pi^{k}(x_{1}))$ (19)
@@
Now, we bound $\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}[f^{k}_{h}(x_{h},a_{h})]$ by $HV^{\pi^{k}}$ plus some lower-order error terms, which we achieve with a ‘self-bounding’ lemma:
Lemma 13.
Define $\delta^{\operatorname{Ber}}_{h}$ that uses $\mathcal{T}^{\pi}$ instead of $\mathcal{T}^{\star}$:
@@
Proof of Lem. 13.
Fix any $f,\pi$. We use the shorthand $\delta_{t}(x,a)=\delta_{t}^{\operatorname{Ber}}(f,\pi,x,a)$ to simplify notation. The corollary follows from the main claim via $\mathbb{E}_{\pi}[Q^{\pi}_{h}(x_{h},a_{h})]\leq V^{\pi}$, since costs are non-negative. To prove the main claim, we establish the following claim by induction:
$\textstyle f_{h}(x_{h},a_{h})\leq\sum_{t=h}^{H}$ $\textstyle(1+\frac{1}{H})^{t-h}\mathbb{E}_{\pi}[c_{t}+$
$\textstyle 28H\delta_{t}(x_{t},a_{t})\mid x_{h},a_{t}].$ (20)
The base case of $h=H+1$ holds since $f_{H+1}=0$. For the induction step, fix any $h\in[H]$ and suppose that Eq. 20 is true for $h+1$. By Eq. 2 and AM-GM, we have
$\textstyle f_{h}(x_{h},a_{h})$ $\leq\textstyle(1+\frac{1}{H})\mathcal{T}^{\pi}_{h}f_{h+1}(x_{h},a_{h})+28H\delta_{h}(x_{h},a_{h})$
By definition, $\mathcal{T}^{\pi}_{h}f_{h+1}(x_{h},a_{h})=\mathbb{E}_{\pi}[c_{h}+f_{h+1}(x_{h+1},a_{h+1})\mid x_{h},a_{h}]$, so we can apply induction hypothesis to $f_{h+1}$. This proves the inductive claim Eq. 20. Then, we prove the main claim by using the fact $(1+\frac{1}{H})^{H}\leq e$. The corollary then follows by $\mathbb{E}_{\pi}[Q^{\pi}_{h}(x_{h},a_{h})]\leq V^{\pi}$ which holds due to the non-negativity of costs. ∎
Thus, by Lem. 13, we can bound Eq. 19 by
@@
By Lem. 12 and the pigeonhole principle, the error terms $\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}$ can be bounded similarly as in the squared loss proof: we can bound $\sum_{k=1}^{K}\delta_{\operatorname{Ber}}^{\operatorname{RL}}(f^{k},\pi^{k})$ by $\widetilde{\mathcal{O}}(d_{\operatorname{bce}}H\beta)$ if UA is false, and by $\widetilde{\mathcal{O}}(Ad_{\operatorname{bce}}^{\operatorname{V}}H\beta)$ if UA is true.
Thus, we have proven Thm. 8 except that the $KV^{\star}$ term is replaced by $\sum_{k=1}^{K}V^{\pi^{k}}$. We show that a first-order bound that scales with $\sum_{k=1}^{K}V^{\pi^{k}}$ implies the seemingly tighter bound that scales with $KV^{\star}$.
Lemma 14.
If $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq c\sqrt{\sum_{k=1}^{K}V^{\pi^{k}}}+c^{2}$, then $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq c\sqrt{2KV^{\star}}+3c^{2}$.
Proof.
By AM-GM, the premise implies $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\frac{1}{2}\sum_{k=1}^{K}V^{\pi^{k}}+\frac{3c^{2}}{2}$, which simplies to $\sum_{k=1}^{K}V^{\pi^{k}}\leq 2KV^{\star}+3c^{2}$. Hence, plugging this back into the premise yields the desired bound. ∎
This concludes the proof of Thm. 8. ∎
4.5 Second-Order Bounds for Online RL with Optimistic MLE: Benefits of Distributional RL
@@
4:     Get optimistic $p^{k}\leftarrow\mathop{\rm arg\,min}_{f\in\mathcal{P}_{k}}\min_{a}\bar{p}_{1}(x_{1},a)$.
@@
To ensure that distributional TD learning succeeds, we assume distributional BC (DistBC) [68].
@@
Then, also define the eluder dimension for mle loss. Let $\Psi^{\operatorname{mle}}_{h}=\{\delta^{\operatorname{dis}}_{h}(\cdot;p):p\in\mathcal{P}\}$ where
@@
Under Assump. 5, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 3 with the mle loss $\ell_{\operatorname{mle}}$ and $\beta=2\ln(H|\mathcal{P}|/\delta)$ enjoys the following:
$\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\textstyle\widetilde{\mathcal{O}}(H\sqrt{{\color[rgb]{1,0,0}\sum_{k=1}^{K}\sigma^{2}(\pi^{k})}\cdot d\beta}+H^{2.5}d\beta),$
@@
The above mle loss bounds scales with the variances of the policies selected by the algorithm, and are thus are called second-order (a.k.a. variance dependent) bounds. As we saw in the CSC setting, a second-order bound is actually strictly sharper than the first-order bound and this is also true in RL [62, Theorem 2.1]. The variance bound can be much tighter in near-deterministic settings where the optimal policy’s cost is far from zero.
While DistRL achieves the tightest scaling w.r.t. $K$ in this paper, it does have one drawback compared to the bce result (Thm. 8): DistRL requires modeling full distributions while the bce loss only requires modeling the mean. Specifically, DistBC is a stronger requirement than the standard BC, and the conditional distribution class $\mathcal{P}$ is in general larger than the regression class $\mathcal{F}$. Nevertheless, in low-rank MDPs with discrete cost distributions, this assumption can always be satisfied with a linear function class of distributions [62, Section 5.1], which also has bounded bracketing entropy of $\widetilde{\mathcal{O}}(dM+\log|\Phi|)$ where $M$ is the number of discretizations. Moreover, distributional TD may not be so restrictive beyond low-rank MDPs, since deep RL approaches with non-linear neural networks have been quite effective for estimating the cost-to-go distributions [9, 16, 10]. Closing this gap between theory and practice is an interesting direction.
We also highlight two important recent works on second-order bounds for contextual bandits, the one-step special case of RL where DistBC simplifies to distributional realizability (Assump. 2). In this setting, Pacchiano [52] proved a second-order bound for contextual bandits with only mean realizability, which is weaker than distributional realizability, by using thresholded least squares. Concurrently, Jia et al. [31] obtained a similar bound as [52] for the strong adversary setting and a complementary bound for the weak adversary setting. Jia et al. [31] also furnished a lower bound proving that the eluder-based second-order bounds in Thm. 9 and [62] are tight and unimprovable even if the number of actions is less than the eluder dimension.
@@
Recall we have that $\delta_{h}^{\operatorname{dis}}\leq\mathcal{E}_{h}^{\operatorname{mle}}$ by Lem. 4. We also write $\delta_{\operatorname{dis}}^{\operatorname{RL}}=\sum_{h=1}^{H}\delta^{\operatorname{dis}}_{h}$ and $\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{mle}}_{h}$. We now establish optimism.
Lemma 15.
Let $\ell=\ell_{\operatorname{mle}}$ and $\mathcal{D}_{h}$ be the same as in Lems. 7 and 12. Then, under Assump. 5, for any $\delta\in(0,1)$ let $\beta=2\ln(H|\mathcal{P}|/\delta)$ and define
@@
W.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}(\hat{p}^{\operatorname{op}},\pi^{i})\leq 2H\beta$ and (b) $\min_{a}\overline{\hat{p}^{\operatorname{op}}_{1}}(x_{1},a)\leq V^{\star}$.
Proof.
The proof follows similarly as the bce case of Lem. 12. By Lem. 3 extended on martingale sequences [3], w.p.a.l. $1-\delta$, for all $p\in\mathcal{P}$, $h\in[H]$,
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{dis}}_{h}(p,\pi^{i})$ $\textstyle\leq\ln(H|\mathcal{P}|/\delta)+\frac{1}{2}L_{h}^{\operatorname{mle}}(p_{h},p_{h+1},\mathcal{D}_{h})$
$\textstyle-\frac{1}{2}L_{h}^{\operatorname{mle}}(\mathcal{T}_{h}^{\operatorname{D},\star}p_{h+1},p_{h+1},\mathcal{D}_{h}).$ (22)
Let $g^{p}_{h}:=\mathop{\rm arg\,min}_{g_{h}\in\mathcal{P}_{h}}L_{h}^{\operatorname{mle}}(g_{h},p_{h+1},\mathcal{D}_{h})$ denote the empirical maximum likelihood estimate. Under the distributional BC premise, we have
$\textstyle\sum_{i=1}^{n}\mathcal{E}^{\operatorname{dis}}_{h}(p,\pi^{i})\leq\ln(H|\mathcal{P}|/\delta)$
$\textstyle+\frac{1}{2}L_{h}^{\operatorname{mle}}(p_{h},p_{h+1},\mathcal{D}_{h})-\frac{1}{2}L_{h}^{\operatorname{mle}}(g^{p}_{h},p_{h+1},\mathcal{D}_{h}).$
Thus, any $p\in\mathcal{C}_{\beta}^{\operatorname{mle}}(\mathcal{D})$ satisfies $\sum_{i=1}^{n}\mathcal{E}^{\operatorname{dis}}_{h}(p,\pi^{i})\leq\frac{1}{2}\beta+\ln(H|\mathcal{P}|/\delta)\leq 2\beta$, which proves Claim (a). For Claim (b), we prove that $Z^{\star}\in\mathcal{C}_{\beta}^{\operatorname{mle}}(\mathcal{D})$. By Eq. 22 and non-negativity of $\mathcal{E}^{\operatorname{dis}}$, we have $L_{h}^{\operatorname{mle}}(\mathcal{T}_{h}^{\operatorname{D},\star}p_{h+1},p_{h+1},\mathcal{D}_{h})-L_{h}^{\operatorname{mle}}(g_{h}^{p},p_{h+1},\mathcal{D}_{h})\leq 2\ln(H|\mathcal{P}|/\delta)=\beta$. Then, setting $p=Z^{\star}$ and noting that $Z_{h}^{\star}=\mathcal{T}_{h}^{\operatorname{D},\star}Z^{\star}_{h+1}$ shows that $Z^{\star}$ satisfies the confidence set condition. Thus, $Z^{\star}\in\mathcal{C}_{\beta}^{\operatorname{mle}}(\mathcal{D})$ and Claim (b) follows by definition of $\hat{p}^{\operatorname{op}}$. ∎
By Lem. 15, we have $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\sum_{k=1}^{K}V^{\pi^{k}}-\min_{a}\bar{p}^{k}_{1}(x_{1},a)$. Now we apply the second-order lemma:
$\textstyle\sum_{k=1}^{K}V^{\pi^{k}}-\bar{p}^{k}_{1}(x_{1},\pi^{k}(x_{1}))$ (23)
@@
Now, we bound the variance term by $\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}[\sigma^{2}(c_{h}+V^{\pi^{k}}_{h+1}(x_{h+1}))]$ plus some lower-order error terms. We achieve this with the following lemma, which can be viewed as a variance analog of Lem. 13.
Lemma 16.
Define the state-action analog of $\delta^{\operatorname{dis}}_{h}$:

### v3

@@ 6.1 First-Order Bounds with BCE Regression
Under Asm. 3, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 2 with the bce loss $\ell_{\operatorname{bce}}$ and $\beta=2\ln(H|\mathcal{F}|/\delta)$ enjoys the following:
$\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\star})\leq\widetilde{\mathcal{O}}(H\sqrt{{\color[rgb]{1,0,0}V^{\star}}K\cdot d\beta}+H^{2}d\beta),$
@@
We remark that Thm. 8 uses the same completeness assumption as Thm. 7, although its bound contains a slightly different eluder dimension $d_{\operatorname{bce}}$ instead of $d_{\operatorname{sq}}$. The eluder dimension for bce is different because the bce loss naturally measures Bellman error via the Bernoulli squared hellinger while the squared loss uses squared distance. The change to $\ell_{1}$ eluder is due to the generalization bound analysis for log-losses, although it is actually sharper than $\ell_{2}$ eluder [44, 64]. However, this is not a significant change for our MDP of interest: Lem. 10 ensures that both eluder dimensions are bounded by $\widetilde{\mathcal{O}}(d)$ in low-rank MDPs with rank $d$. Indeed, this implies that the first-order bound in Thm. 8 can be specialized for low-rank MDPs by the same argument as before, yielding the first small-loss bound for low-rank MDPs in online RL without distributional RL [64]. Characterizing the exact differences between these eluder dimension variants is an interesting question for future work.
In terms of related works, several other works also propose to use different losses than squared loss to evaluate the Bellman error. For example, Bas-Serrano et al. [9] argued that the logistic loss is advantageous since it is convex, whereas the squared loss is not convex in the $Q$-function due to the $\max$ operator. Farebrother et al. [21] is a more applied paper which shows that classification losses such as cross-entropy scale much better with deep networks than squared loss. Overall these other works provide other reasons for why squared loss is sub-optimal, while we focus on the improved sample efficiency aspect of employing alternative losses. We now prove Thm. 8.
@@
Note that $\mathcal{T}_{h}^{\star}f_{h+1}(x_{h},a_{h})=\mathbb{E}[\tau^{\star}(f_{h+1},c_{h},x_{h+1})\mid x_{h},a_{h}]$, which is realizable by BC. Recall that $\delta_{h}^{\operatorname{Ber}}\leq\mathcal{E}_{h}^{\operatorname{bce}}$ by Lem. 2. We also write $\delta_{\operatorname{Ber}}^{\operatorname{RL}}=\sum_{h=1}^{H}\delta^{\operatorname{Ber}}_{h}$ and $\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{bce}}_{h}$. The following lemma establishes optimism and is analogous to Sec. 5. {restatable}lemmaBCERLOptimism Let $\ell=\ell_{\operatorname{bce}}$. Under the same setup as Sec. 5 with $\hat{f}^{\operatorname{op}}$ selected from $\mathcal{C}^{\operatorname{bce}}_{\beta}$ instead of $\mathcal{C}^{\operatorname{sq}}_{\beta}$, w.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}(\hat{f}^{\operatorname{op}},\pi^{i})\leq 2H\beta$, and (b) $\min_{a}\hat{f}^{\operatorname{op}}_{1}(x_{1},a)\leq V^{\star}$. By Sec. 6.1, we have $\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\star})\leq\sum_{k=1}^{K}(V^{\pi^{k}}-\min_{a}f^{k}_{1}(x_{1},a))$. Then, the proof follows similarly as the squared loss case from before, except that we apply the finer Eq. 2 in place of Cauchy-Schwarz:
$\textstyle\sum_{k=1}^{K}(V^{\pi^{k}}-f^{k}_{1}(x_{1},\pi^{k}(x_{1})))$ (17)
@@
Now, we bound $\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}[f^{k}_{h}(x_{h},a_{h})]$ by $HV^{\pi^{k}}$ plus some lower-order error terms, which we achieve with a ‘self-bounding’ lemma: {restatable}lemmaSelfBoundingBCELemma Define $\delta^{\operatorname{Ber}}_{h}$ that uses $\mathcal{T}^{\pi}$ instead of $\mathcal{T}^{\star}$:
@@
By Sec. 6.1, we can bound Eq. 17 by
@@
By Sec. 6.1 and the pigeonhole principle, the error terms $\mathcal{E}_{\operatorname{bce}}^{\operatorname{RL}}$ can be bounded similarly as in the squared loss proof: we can bound $\sum_{k=1}^{K}\delta_{\operatorname{Ber}}^{\operatorname{RL}}(f^{k},\pi^{k})$ by $\widetilde{\mathcal{O}}(d_{\operatorname{bce}}H\beta)$ if UA is false, and by $\widetilde{\mathcal{O}}(Ad_{\operatorname{bce}}^{\operatorname{V}}H\beta)$ if UA is true.
Thus, setting $d=d_{\operatorname{bce}}$ if UA is false, and $d=Ad_{\operatorname{bce}}^{\operatorname{V}}$ if UA is true, we have proven
$\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\star})\leq H\sqrt{\sum_{k=1}^{K}V^{\pi^{k}}\cdot d\beta}+H^{2}d\beta.$
Finally, we observe an implicit inequality where we can replace $\sum_{k=1}^{K}V^{\pi^{k}}$ by $KV^{\star}$ by collecting a factor of $3$, as shown by the following lemma. {restatable}lemmaImplicitFirstOrderIneq If $\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\star})\leq c\sqrt{\sum_{k=1}^{K}V^{\pi^{k}}}+c^{2}$, then $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq c\sqrt{2KV^{\star}}+3c^{2}$. This concludes the proof of Thm. 8. ∎
6.2 Second-Order Bounds with MLE
@@
4:     Get optimistic $p^{k}\leftarrow\mathop{\rm arg\,min}_{p\in\mathcal{P}_{k}}\min_{a}\bar{p}_{1}(x_{1},a)$.
@@
To ensure that distributional TD learning succeeds, we assume distributional BC (DistBC) [72, 64].
@@
Distributional BC posits that the distribution class is closed under the distributional Bellman operator and is a stronger condition than the standard completeness condition (Asm. 3). Nevertheless, in low-rank MDPs with discrete cost distributions, Asm. 5 can be satisfied with a linear distribution class [65, Section 5.1], which also has bounded bracketing entropy of $\widetilde{\mathcal{O}}(dM+\log|\Phi|)$ where $M$ is the number of discretizations.
Next, we define the eluder dimension for mle loss. Let $\Psi^{\operatorname{mle}}_{h}=\{\delta^{\operatorname{dis}}_{h}(\cdot;p):p\in\mathcal{P}\}$ where
@@
Under Asm. 5, for any $\delta\in(0,1)$, w.p.a.l. $1-\delta$, Alg. 3 with the mle loss $\ell_{\operatorname{mle}}$ and $\beta=2\ln(H|\mathcal{P}|/\delta)$ enjoys the following:
$\sum_{k=1}^{K}(V^{\pi^{k}}-V^{\star})\leq\textstyle\widetilde{\mathcal{O}}(H\sqrt{{\color[rgb]{1,0,0}\sum_{k=1}^{K}\sigma^{2}(\pi^{k})}\cdot d\beta}+H^{2.5}d\beta),$
@@
The above mle loss bounds scales with the variances of the policies selected by the algorithm, and are thus are called second-order (a.k.a. variance dependent) bounds. As we saw in the CSC setting, a second-order bound is actually strictly sharper than the first-order bound and this is also true in RL [65, Theorem 2.1]. The variance bound can be much tighter in near-deterministic settings where the optimal policy’s cost is far from zero.
One drawback of the DistRL approach is that it requires modeling the entire conditional distributions which are more complex than the conditional mean, i.e., $\mathcal{P}$ is generally larger and more complex than $\mathcal{F}$ from Thms. 7 and 8. Also, DistRL requires completeness w.r.t. the distributional Bellman operator, which is stronger than standard BC. However, it is worth noting that in practice DistRL is often much more performant than non-distributional approaches, which suggests that these modeling conditions (e.g., distributional completeness) are not so restrictive [10, 17, 21]. Closing this gap between theory and practice is an interesting future direction.
We also highlight two recent works on second-order bounds for contextual bandits, the one-step special case of RL where DistBC simplifies to distributional realizability (Asm. 2). In this setting, Pacchiano [55] proved a second-order bound for contextual bandits with only mean realizability, which is weaker than distributional realizability, by using thresholded least squares. Concurrently, Jia et al. [33] obtained a similar bound as [55] for the strong adversary setting and a complementary bound for the weak adversary setting. Jia et al. [33] also furnished a lower bound proving that the eluder-based second-order bounds in Thm. 9 and [65] are tight and unimprovable even if the number of actions is less than the eluder dimension.
Computational considerations for bce and mle. We remark that switching from squared loss to bce does not incur any computational overhead and is a single line of code in practice. The mle loss is used for distribution fitting and thus the function approximator should output a distribution instead of a scalar. In practice, the distribution can be modeled with histograms [10, 31, 21] or quantiles [17] and is only a constant factor more to compute and maintain than the non-distributional losses. Thus, the mle loss and distributional RL more generally also do not incur any notable computational overhead in practice.
@@
Recall we have that $\delta_{h}^{\operatorname{dis}}\leq\mathcal{E}_{h}^{\operatorname{mle}}$ by Lem. 4. We also write $\delta_{\operatorname{dis}}^{\operatorname{RL}}=\sum_{h=1}^{H}\delta^{\operatorname{dis}}_{h}$ and $\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}=\sum_{h=1}^{H}\mathcal{E}^{\operatorname{mle}}_{h}$. We now establish the optimism lemma, which is analogous to Sec. 5 and Sec. 6.1. {restatable}lemmaMLERLOptimism Let $\ell=\ell_{\operatorname{mle}}$ and $\mathcal{D}_{h}$ be the same as in Secs. 5 and 6.1. Then, under Asm. 5, for any $\delta\in(0,1)$ let $\beta=2\ln(H|\mathcal{P}|/\delta)$ and define
@@
W.p.a.l. $1-\delta$, we have (a) $\sum_{i=1}^{n}\mathcal{E}_{\operatorname{mle}}^{\operatorname{RL}}(\hat{p}^{\operatorname{op}},\pi^{i})\leq 2H\beta$ and (b) $\min_{a}\overline{\hat{p}^{\operatorname{op}}_{1}}(x_{1},a)\leq V^{\star}$. By Sec. 6.2, we have $\sum_{k=1}^{K}V^{\pi^{k}}-V^{\star}\leq\sum_{k=1}^{K}V^{\pi^{k}}-\min_{a}\bar{p}^{k}_{1}(x_{1},a)$. Now we apply the second-order lemma:
$\textstyle\sum_{k=1}^{K}(V^{\pi^{k}}-\bar{p}^{k}_{1}(x_{1},\pi^{k}(x_{1})))$ (19)
@@
Now, we bound the variance term by $\sum_{h=1}^{H}\mathbb{E}_{\pi^{k}}[\sigma^{2}(c_{h}+V^{\pi^{k}}_{h+1}(x_{h+1}))]$ plus some lower-order error terms. We achieve this with the following lemma, which can be viewed as a variance analog of Sec. 6.1. {restatable}lemmaVarianceChangeOfMeasure Define the state-action analog of $\delta^{\operatorname{dis}}_{h}$: